### v1

# inGRASS: Incremental Graph Spectral Sparsification via Low-Resistance-Diameter Decomposition

Abstract
This work presents inGRASS, a novel algorithm designed for incremental spectral sparsification of large undirected graphs. The proposed inGRASS algorithm is highly scalable and parallel-friendly, having a nearly-linear time complexity for the setup phase and the ability to update the spectral sparsifier in $O(\log N)$ time for each incremental change made to the original graph with $N$ nodes. A key component in the setup phase of inGRASS is a multilevel resistance embedding framework introduced for efficiently identifying spectrally-critical edges and effectively detecting redundant ones, which is achieved by decomposing the initial sparsifier into many node clusters with bounded effective-resistance diameters leveraging a low-resistance-diameter decomposition (LRD) scheme. The update phase of inGRASS exploits low-dimensional node embedding vectors for efficiently estimating the importance and uniqueness of each newly added edge. As demonstrated through extensive experiments, inGRASS achieves up to over 200×200$\times$200 × speedups while retaining comparable solution quality in incremental spectral sparsification of graphs obtained from various datasets, such as circuit simulations, finite element analysis, and social networks.
I Introduction
The graph-based analysis is a crucial technique that finds extensive application in various electronic design automation (EDA) problems like logic synthesis and verification, layout optimization, static timing analysis (STA), network partitioning/decomposition, circuit modeling, and simulation. For instance, spectrally-sparsified graphs allow for accelerating circuit simulations, performing vectorless integrity verification of power grids, and analyzing worst-case on-chip temperature distributions.
In recent years, mathematics and theoretical computer science (TCS) researchers have extensively studied various research problems related to simplifying large graphs using spectral graph theory [1, 2, 3]. Specifically, recent research on spectral graph sparsification allows for the construction of much sparser subgraphs that can preserve important graph spectral (structural) properties like the first few eigenvalues and eigenvectors of the graph Laplacian. These findings have already led to the development of numerical and graph algorithms that can solve large sparse matrices and partial differential equations (PDEs) in nearly-linear time, as well as enable graph-based semi-supervised learning (SSL), computing the stationary distributions of Markov chains and personalized PageRank vectors, spectral graph partitioning and data clustering, max flow and multi-commodity flow of undirected graphs, and nearly-linear time circuit simulation and verification algorithms [4, 2, 5, 6].
However, there still remain grand challenges when adopting spectral sparsification algorithms to real-world EDA applications: existing spectral sparsification methods can not efficiently update the sparsified graph when only incremental changes are made to the original input graph [5, 7, 8, 9, 10]. For example, when a power grid network has been updated with a few additional metal wires connected to the system, traditional spectral sparsification methods (e.g. GRASS [7], feGRASS [8]) must recompute the sparsifier from scratch, imposing a significant overhead during the chip optimization procedure. While dynamic algorithms for spectral graph sparsification have been recently studied to handle streaming edge insertions/deletions [11, 12], it remains unclear if such theoretical results would allow for practically efficient implementations.
Figure 1: The proposed inGRASS algorithm for incremental spectral sparsification. Given the initial input graph $G^{(0)}$ and its sparsifier $H^{(0)}$, inGRASS constructs the updated sparsifiers $H^{(1)},H^{(2)},\cdots$ with newly added edges.
To address the aforementioned limitations, this work presents a highly-efficient algorithmic framework (inGRASS) for incremental spectral graph sparsification (as shown in Fig. 1). A key component of inGRASS is an efficient effective-resistance embedding scheme in the setup phase, which leverages low-resistance-diameter (LRD) decomposition to partition the original sparsifier into multiple node clusters with bounded resistance diameters. We show that by exploiting a multilevel LRD decomposition procedure, each node can be represented by a low-dimensional embedding vector that allows for extremely fast estimation of effective-resistance distances between any two nodes in the sparsifier. Motivated by recent spectral perturbation-based spectral sparsification methods [5, 7, 9] that prune each off-tree edge candidate based on its spectral distortions (defined as the product of the edge weight and effective resistance in the sparsifier), in the incremental update phase inGRASS handles each newly added edge by checking its spectral distortion that can be quickly estimated based on the proposed resistance embedding scheme, which allows for identifying the most spectrally-critical edges while filtering out non-critical ones in the current sparsifier. Compared with re-running spectral sparsification algorithms from scratch, inGRASS has a much lower computational complexity for incremental edge insertions: for a weighted undirected graph with $O(N)$ nodes, the setup phase of inGRASS enjoys nearly-linear $O(N\log N)$ time complexity as prior methods, while each incremental update of the sparsified graph can be accomplished in $O(\log N)$ time. The key contribution of this work is summarized below:
  1. 1.
    We propose an incremental spectral graph sparsification (inGRASS) algorithm for efficiently updating sparsified graphs considering newly added edges.
  2. 2.
    A key component of inGRASS is an effective-resistance embedding scheme that leverages multilevel LRD decomposition and allows for extremely fast effective-resistance calculations critical to spectral distortion analysis of each new edge inserted into the original graph.
  3. 3.
    Our extensive experiment results on real-world VLSI designs show that inGRASS is over 200×200$\times$200 × faster than running state-of-the-art spectral sparsification algorithms from scratch while achieving comparable solution quality (condition number, graph density, etc).
The rest of the paper is organized as follows. In Section II, we provide a background introduction to the basic concepts related to spectral graph sparsification. In section III, we introduce the technical details of inGRASS as well as its algorithm flow and complexity. In Section IV, we demonstrate extensive experimental results to evaluate the performance of inGRASS using a variety of real-world VLSI design benchmarks, which is followed by the conclusion of this work in Section V.
II Background
II-A Spectral Graph Theory
In a weighted undirected graph $G=(V,E,w)$, $V$ ($|V|=N$) and $E$ denote the sets of vertices and edges, $w$ is a positive weight function, and $w_{i,j}$ denotes the weight of the edge between vertices $i$ and $j$, or $w(e)$ for edge $e\in E$. The $N\times N$ adjacency matrix $A$ is symmetric ($A(i,j)=A(j,i)$) and positive semi-definite ($x^{\top}Ax\geq 0$ for any real vector $x\in\mathcal{R}^{N\times 1}$). The Laplacian matrix $L_{G}$ is defined as $L_{G}:=D-A$, where $D$ is the degree matrix. Eigenvalues and eigenvectors of $L_{G}$ reveal properties related to the graph structure, such as the number of connected components and node connectivity. The Laplacian quadratic form, $x^{\top}L_{G}x$, is utilized in spectral graph theory to assess properties like graph cuts, clustering, and conductance, serving as an analytical tool for characterizing the Laplacian matrix $L_{G}$ and the associated graph structure. It is also suggested that the spectral similarity between two graphs $G$ and $H$ can be measured using the following inequality [13]:
$\frac{x^{\top}L_{G}x}{\epsilon}\leq x^{\top}L_{H}x\leq\epsilon x^{\top}L_{G}x,$ (1)
where $L_{G}$ and $L_{H}$ are the Laplacian matrices of $G$ and $H$, respectively. Similarly, a smaller relative condition number $\kappa(L_{G},L_{H})$ implies a higher degree of spectral similarity between $G$ and $H$.
II-B Spectral Graph Decomposition
Lemma II.1
Spectral sparsification of an undirected graph $G=(V,E,w)$, with its Laplacian denoted by $L_{G}$, can be achieved by leveraging a short-cycle decomposition algorithm. This algorithm produces a sparsified graph $H=(V,E^{\prime},w^{\prime})$, where $E^{\prime}<<E$, with its Laplacian denoted by $L_{H}$, such that for all real vectors $x\in\mathcal{R}^{V}$, the approximation $x^{\top}L_{G}x\approx x^{\top}L_{H}x$ holds [14].
Building upon Lemma II.1, the graph sparsification algorithm combines short-cycle decomposition with low-stretch spanning trees (LSSTs) [15] to construct a sparsified graph that preserves the spectral properties of the original graph [14].
Furthermore, a similar spectral graph sparsification method is proposed in [9], which decomposes the graph into multiple sets of disjoint cycles using a multilevel spectral graph coarsening framework. By constructing a hierarchy of contracted graphs, the method identifies and adds spectrally-critical edges to the initial sparse graph.
II-C Effective Resistance
For a weighted, undirected graph $G=(V,E,w)$ with $|V|=N$, the effective resistance between nodes $(p,q)\in V$ plays a crucial role in various graph analysis tasks including spectral sparsification algorithms [13]. The effective resistance distances can be accurately computed using the equation:
$R_{eff}(p,q)=\sum\limits_{i=2}^{N}\frac{w_{p,q}(u_{i}^{\top}b_{pq})^{2}}{\lambda_{i}},$ (2)
where ${b_{p}}\in\mathbb{R}^{V}$ denote the standard basis vector with all zero entries except for the $p$-th entry being $1$, and $q$-th entry being $-1$ (${b_{pq}}=b_{p}-b_{q}$). $u_{i}\in\mathbb{R}^{\mathcal{V}}$ for $i=1,...,|V|$ denote the unit-length, mutually-orthogonal eigenvectors corresponding to Laplacian eigenvalues $\lambda_{i}$ for $i=1,...,|V$.
III inGRASS: Incremental Graph Sparsification
In this section, we introduce inGRASS, an incremental graph spectral sparsification method in response to a stream of edge insertions, achieving nearly-linear time complexity. The key strength of inGRASS lies in its ability to rapidly identify spectrally-critical edges within the stream without the need for recomputing the edge importance metric for each new arrival. This feature significantly reduces the computational complexity of the algorithm, making inGRASS outperform existing graph sparsification and incremental edge update methods.
III-A Overview of inGRASS Algorithm
The proposed inGRASS algorithm constructs a sparse data structure to enable fast estimation of the spectral distortion and spectral similarity of newly introduced edges. It consists of two primary phases: setup phase and update phase.
Setup Phase
This is a one-time operation that equips the inGRASS algorithm to iteratively update the initial sparse graph $H$ upon receiving streams of newly introduced edges in different iterations. In this phase, each node in the initial sparse graph $H^{(0)}$ is assigned a $\log N$-dimensional embedding vector by leveraging low-resistance-diameter (LRD) decomposition. A sparse data structure is then created that provides an efficient estimate of the resistance distance among different pair of nodes in the initial sparse graph $H^{(0)}$.
Update Phase
Computing $O(\log N)$ node embedding vectors in the previous phase enables efficient spectral distortion estimation for newly added edges, identifying spectrally-critical edges. Subsequently, spectral similarity estimation is employed to enhance sparsification performance by selectively filtering edges with minimal impact on graph spectral properties. The determination of the filtering level is based on the target relative condition number and can be adjusted to achieve various degrees of spectral similarity between graph $G$ and graph $H$.
Combining these setup and update phases, inGRASS presents an efficient approach for incremental graph spectral sparsification, outperforming the existing methods.
III-B The Setup Phase of inGRASS
III-B1 Scalable Estimation of Effective Resistances
To address the computational complexity associated with directly computing eigenvalues and eigenvectors required for estimating edge effective resistances, we introduce a scalable algorithm that approximates the eigenvectors of the graph Laplacian matrix using the Krylov subspace.
Figure 2: A 4-level resistance embedding of the initial sparse graph achieved through the proposed LRD decomposition. Since the embedding vectors for nodes $5$ and $9$ are $[3,3,3,2]^{\top}$ and $[6,2,2,2]^{\top}$, the effective-resistance distance between them is bounded by the resistance diameter of cluster $2$ shown in (d).
Let $A_{N\times N}$ denote the adjacency matrix of a graph, its Krylov subspace $\mathbf{K}_{m}(A,x)$ is a vector space spanned by the vectors computed through power iterations $x,Ax,A^{2}x,\ldots,A^{m-1}x$ [16]. By enforcing orthogonality among the above vectors in the Krylov subspace, we can compute a new set of mutually-orthogonal vectors of unit lengths for approximating the original Laplacian eigenvectors in (2), which are denoted as $\tilde{u}_{1}$, $\tilde{u}_{2}$, $\ldots$, $\tilde{u}_{m}$. To estimate the effective resistance between two nodes $p$ and $q$, we leverage the approximated eigenvectors of the Laplacian matrix $L_{G}$ constructed through the above Krylov subspace. Using (2) we obtain an estimation of each edge’s effective resistance:
$R_{eff}(p,q)\approx\sum\limits_{i=1}^{m}\frac{w_{p,q}(\tilde{u}^{\top}_{i}b_{pq})^{2}}{\tilde{u}^{\top}_{i}L_{G}\tilde{u}_{i}},$ (3)
where $\tilde{u}_{i}$ represents the approximated eigenvector corresponding to the $i$-th eigenvalue of $L_{G}$, and $m$ is the order of the Krylov subspace.
III-B2 Multilevel Resistance Embedding
We utilize a multilevel low-resistance-diameter (LRD) decomposition framework for embedding nodes of the initial sparse graph $H^{(0)}$ into an $O(\log N)$-dimensional space [17]. The embedding process iteratively follows these steps: (S1) efficiently estimate the effective resistance of edges using (3); (S2) contract edges starting from low-resistance-diameter clusters (cluster diameters are initialized to 0 for all nodes in the first level); (S3) replace a contracted edge with a supernode and adjust the resistance diameter of the clusters accordingly.
III-B3 Multilevel Sparse Data Structure
Leveraging the $O(\log N)$-dimensional node embedding vectors obtained in the previous step, we construct a specialized sparse data structure for efficient access to the node clustering indices at different levels. This facilitates efficient storage and retrieval of cluster connectivity information, enabling the identification of spectrally-critical and spectrally-unique edges among the newly added edges. The sparse data structure is promptly updated upon the addition of a newly introduced edge to the sparse graph.
Figure 2 illustrates the computation of node embedding vectors through multilevel LRD decomposition of the initial sparse graph. Increasing the diameter thresholds, as shown in the figure, leads to larger cluster sizes progressing from left to right. This allows for estimating the upper bound of effective resistance between any two nodes based on the level where they share the same cluster index. For instance, nodes 5 and 9, with different cluster indices at levels (a), (b), and (c), share the same cluster index at level (d). Their resistance distance is bounded by the resistance diameter of cluster 2 at level (d).
III-C The Update Phase of inGRASS
III-C1 Spectral Distortion Estimation
It has been demonstrated that inserting a new edge with a large spectral distortion  [^1] The spectral distortion of an edge is defined as the product of its edge weight and the effective resistance between its two end nodes [7, 9]. to the sparse graph will significantly increase its first few Laplacian eigenvalues [9, 7].
Lemma III.1
Let $H=(V,E^{\prime},w^{\prime})$, where $w^{\prime}:E^{\prime}\rightarrow\mathbb{R}+$, denote the sparsified weighted graph of $G$, and $L_{H}$ denote its Laplacian matrix. The $i$-th Laplacian eigenvalue perturbation due to $\delta L_{H}=w^{\prime}_{p,q}b_{pq}b^{\top}_{pq}$ can be computed as:
$\delta\lambda_{i}=w^{\prime}_{p,q}(u_{i}^{\top}b_{pq})^{2},$ (4)
where $u_{i}$ represents the eigenvector corresponding to the $i$-th eigenvalue $\lambda_{i}$ of the Laplacian matrix $L_{H}$.
Lemma III.2
Construct a weighted eigensubspace matrix $U_{K}$ for $K$-dimensional spectral graph embedding using the first $K$ Laplacian eigenvectors and eigenvalues as follows:
$U_{K}=\left[\frac{u_{2}}{\sqrt{\lambda_{2}}},...,\frac{u_{K}}{\sqrt{\lambda_{K}}}\right],$ (5)
then the spectral distortion of the new edge $e_{p,q}$ will become the total $K$-eigenvalue perturbation $\Delta_{K}$ when $K\rightarrow N$ [9]:
$\Delta_{K}=\sum\limits_{i=2}^{{K}}\frac{\delta{{\lambda}_{i}}}{\lambda_{i}}=w_{p,q}\|U_{K}^{\top}b_{pq}\|^{2}_{2}\approx w_{p,q}R(p,q).$ (6)
The equation (6) proves that a higher spectral distortion $\Delta_{K}$ can be achieved by including edges with higher effective resistance $R(p,q)$ and higher edge weight $w_{p,q}$. Our algorithm efficiently estimates the effective resistance of the newly introduced edges, leveraging the node embedding information obtained from the setup phase. This enables sorting the newly introduced edges according to their estimated spectral distortion that leads to including the spectrally-critical edges to the sparse graph.
III-C2 Spectral Similarity Estimation
The inGRASS algorithm applies an edge filtering process to exclude a newly added edge if there is already an existing edge in the sparse graph with a similar spectral distortion. This filtering process minimizes the introduction of additional edges while still influencing the spectral distortion. The algorithm selects a filtering level ($\mathcal{L}$) from the levels computed during the LRD decomposition phase with respect to a target condition number. For a target condition number $\kappa(L_{G},L_{H})=C$, the filtering level ($\mathcal{L}$) is chosen with the maximum number of nodes in a cluster equal to $\frac{C}{2}$. This threshold bounds the maximum spectral distortion among nodes within clusters (equivalent to the shortest path among the nodes within each cluster) to achieve the desired condition number.
Figure 3: (a) The original graph featuring three newly introduced edges highlighted in red. (b) The edge included into the sparse graph, marked in red, alongside the edges with adjusted weights, denoted in blue.
For a newly introduced edge, inGRASS only includes the edge if there is no existing edge in the sparse graph connecting the clusters to which those two nodes belong in the specified filtering level ($\mathcal{L}$). In the event of an existing edge connecting the clusters of those nodes in the filtering level ($\mathcal{L}$), the newly introduced edge is discarded, and the weight of the existing edge is adjusted by adding the weight of the newly added edge to it. Moreover, if the newly introduced edge connects two nodes within the same cluster in the filtering level ($\mathcal{L}$), the edge is discarded and its weight is proportionally distributed among the edges connecting the nodes within that cluster.
Fig. 3 illustrates the proposed edge filtering process employed by the inGRASS algorithm to select spectrally-critical and unique edges that will significantly impact the global structure of the sparsified graph. Let’s consider three newly added edges to the original graph: $e_{1}=(2,11)$, $e_{2}=(7,13)$, and $e_{3}=(10,14)$, as depicted in Fig. 3 (a), the algorithm determines the filtering level, denoted as $\mathcal{L}$ = (b) in Fig. 2, to achieve a desired condition number of $\kappa(L_{G},L_{H})=8$. Starting with $e_{1}=(2,11)$, the algorithm integrates it into the existing edge $e=(1,2)$ between clusters 2 and 11, adjusting the weight accordingly. For $e_{2}=(7,13)$, both nodes 7 and 13 share a cluster, leading to the exclusion of $e_{2}=(7,13)$ and a proportional increase in weights of related edges. Finally, $e_{3}=(10,14)$ is added as no existing edge connects the clusters of nodes 10 and 14 in the filtering level $\mathcal{L}=$ (b).
III-D The Algorithm Flow and Complexity Analysis
The complete flow of the inGRASS algorithm is described in Algorithm 1 to efficiently update the sparse graph $H^{(0)}=(V,E^{\prime},w^{\prime})$ with $|V|=N$ after adding a set of weighted edges to the original graph $G^{(0)}$ to reach the target condition number $\kappa(L_{G},L_{H})=C$.
Algorithm 1 inGRASS algorithm flow
Input: $H^{(0)}=(V,E^{\prime})$ where $|V|=N$, a set of newly introduced edges to $G^{(0)}$, a target condition number $\kappa(L_{G},L_{H})=C$.
Output: The updated sparse graph $H$.
1:  Setup Phase 1: Effective resistance estimation using (3).
2:  Setup Phase 2: Node embeddings via LRD decomposition.
3:  Setup Phase 3: $O(\log N)$-level sparse data structure.
4:  Update Phase 1: Spectral distortion estimation of new edges.
5:  Update Phase 2: Spectral similarity estimation according to $C$.
6:  Return $H$.
The complexity analysis of the inGRASS algorithm is split into two phases:
  • In the setup phase, the $O(\log N)$-level multilevel low-resistance-diameter (LRD) decomposition is applied to compute node cluster indices, resulting in a complexity of $O(N\log N)$.
  • In the update phase, the complexity arises from calculating the spectral distortion and checking spectral similarity of newly added edges using the $O(\log N)$-dimensional node embedding vectors from the setup phase, with a complexity of $O(\log N)$ for each new edge in the original graph.
IV Experimental results
TABLE I: GRASS time vs inGRASS setup time.
Test Cases $|V|$ $|E|$ GRASS (s) Setup (s)
G3 circuit 1.5E+6 3.0E+6 18.7 s 13.7 s
G2 circuit 1.5E+5 2.9E+5 0.75 s 0.9 s
fe_4elt2 1.1E+4 3.3E+4 0.053 s 0.06 s
fe_ocean 1.4E+5 4.1E+5 1.12 s 1.01 s
fe_sphere 1.6E+4 4.9E+4 0.08 s 0.17 s
delaunay_n18 2.6E+5 6.5E+5 2.2 s 1.9 s
delaunay_n19 5.2E+5 1.6E+6 6.2 s 4 s
delaunay_n20 1.0E+6 3.1E+6 14.1 s 9.5 s
delaunay_n21 2.1E+6 6.3E+6 28.5 s 19 s
delaunay_n22 4.2E+6 1.3E+7 62 s 38.6 s
M6 3.5E+6 1.1E+7 83 s 45 s
 [^3] SP 3.7E+6 1.1E+7 84 s 46 s
AS365 3.8E+6 1.1E+7 84 s 48 s
NACA15 1.0E+6 3.1E+6 13.8 s 8 s
This section presents the results of a diverse range of experiments conducted to assess the performance and efficiency of the proposed incremental graph spectral sparsification algorithm (inGRASS). The test cases were selected from a wide array of matrices commonly utilized in circuit simulation, and finite element analysis applications  [^2] https://sparse.tamu.edu/. All experiments were conducted on a Linux Ubuntu system with 1 terabyte of RAM and a $3.6$ GHz $64$-core CPU. The state-of-the-art spectral sparsification tool GRASS [7]  [^3] https://sites.google.com/mtu.edu/zhuofeng-graphspar/home has been used as the benchmark to evaluate the performance and scalability of inGRASS.
Table I provides a runtime comparison between GRASS and inGRASS setup times (measured in seconds) for various test cases. The runtime for the setup phase of inGRASS is mainly due to the LRD decomposition step for computing resistance embeddings for the initial sparse graph $H^{(0)}$. Note that this step is a one-time task, which can be leveraged for many subsequent update iterations involving streaming edge modifications. We observe that the inGRASS setup time is even faster than the GRASS time for most test cases as shown in Table I, where $|V|$ ($|E|$) denotes the number of nodes (edges) in the original graph $G^{(0)}$, respectively.
IV-A inGRASS for Incremental Spectral Sparsification
TABLE II: Comparison of Incremental Graph Spectral Sparsification Outcomes through $10$-Iterative Updates using GRASS, inGRASS, and Random Methods.
Test Cases Density (D) $\kappa(L_{G},L_{H})$ GRASS-D inGRASS-D Random-D GRASS-T inGRASS-T $\frac{GRASS-T}{inGRASS-T}$
G3 circuit 10.0% $\rightarrow$ 34% 88 $\rightarrow$ 353 11.6% 11.7% 23.0% 196 s 1.7 s 115 $\times$
G2 circuit 9.0% $\rightarrow$ 32% 72 $\rightarrow$ 283 11.0% 11.4% 25.7% 7.8 s 0.11 s 71 $\times$
fe_4elt2 10.0% $\rightarrow$ 39% 95 $\rightarrow$ 330 10.0% 10.1% 36.0% 0.56 s 0.008 s 70 $\times$
fe_ocean 9.8% $\rightarrow$ 50% 210 $\rightarrow$ 468 9.6% 11.5% 30.3% 11.8 s 0.13 s 91 $\times$
fe_sphere 10.5% $\rightarrow$ 41% 123 $\rightarrow$ 1103 8.2% 10.5% 34.9% 0.84 s 0.009 s 93 $\times$
delaunay_n18 10.5% $\rightarrow$ 35% 113 $\rightarrow$ 336 11.7% 11.6% 30.5% 23.1 s 0.19 s 122 $\times$
delaunay_n19 10.6% $\rightarrow$ 35% 122 $\rightarrow$ 406 11.9% 11.8% 29.7% 65.1 s 0.41 s 159 $\times$
delaunay_n20 10.5% $\rightarrow$ 35% 126 $\rightarrow$ 418 11.7% 11.7% 30.2% 148 s 0.9 s 164 $\times$
delaunay_n21 10.0% $\rightarrow$ 35% 151 $\rightarrow$ 428 11.0% 10.9% 29.5% 299 s 2.1 s 142 $\times$
delaunay_n22 10.3% $\rightarrow$ 34% 150 $\rightarrow$ 491 10.7% 10.7% 30.7% 651 s 4.3 s 151 $\times$
M6 9.8% $\rightarrow$ 34% 172 $\rightarrow$ 817 10.2% 11.2% 29.3% 871 s 4 s 218 $\times$
 [^3] SP 9.7% $\rightarrow$ 34% 180 $\rightarrow$ 897 9.5% 11.0% 29.4% 882 s 4.2 s 210 $\times$
AS365 10.1% $\rightarrow$ 34% 157 $\rightarrow$ 1876 10.4% 12.7% 31.1% 885 s 4.5 s 197 $\times$
NACA15 10.4% $\rightarrow$ 34% 152 $\rightarrow$ 585 10.4% 11.5% 29.3% 145 s 1 s 145 $\times$
Table II presents a thorough comparison of our incremental graph spectral sparsification algorithm, inGRASS, with GRASS and Random methods through $10$-iterative updates. The density ($D$), defined as $D:=\frac{|E|}{|V|}$, shows the density of the initial sparse graph $H^{(0)}$ and its density when all newly introduced edges are included. This emphasizes the importance of incremental spectral sparsification to prevent a substantial increase in sparse graph density. We maintain an initial density of $D=10\%$ for consistency. The relative condition number $\kappa(L_{G},L_{H})$ measures the spectral similarity between the original graph $G$ and the updated sparse graph $H$. A smaller relative condition number indicates higher spectral similarity. The column related to $\kappa(L_{G},L_{H})$ indicates how the condition number between $G^{(0)}$ and $H^{(0)}$ is perturbed when newly added edges are excluded from $H^{(0)}$, providing insights into edge generation. Columns ”GRASS-D,” ”inGRASS-D,” and ”Random-D” compare sparse graph density for a target condition number using each method. The target condition number is set to match the initial condition number between $G^{(0)}$ and $H^{(0)}$. For instance, the target condition number for ”G3 circuit” is set to $\kappa(L_{G^{(0)}},L_{H^{(0)}})=88$. Notably, inGRASS achieves the target condition number comparably to GRASS and significantly outperforms Random, accompanied by a runtime speedup exceeding $\frac{GRASS-T}{inGRASS-T}=200\times$ across all test cases through $10$-iterative updates.
IV-B Robustness of the inGRASS Algorithm
The performance of our incremental graph spectral sparsification algorithm, inGRASS, is studied for the ”G2 circuit” dataset over different initial sparse graph densities. In Table III, the labeled column ”Density (D)” shows an initial density range of $6.5\%$ to $12.7\%$, with the sparse graph density set to $32\%$ when all edges are included. The $\kappa(L_{G},L_{H})$ column indicates the perturbation in the condition number between the original graph $G^{(0)}$ and the initial sparse graph $H^{(0)}$ when no newly added edges are included. The target condition number is set to the initial value (e.g. 56 for $D=12.7\%$). Both ”GRASS-D” and ”inGRASS-D” denote the sparse graph density for the same target condition number, with inGRASS showing comparable results to GRASS across various initial densities.
TABLE III: GRASS vs inGRASS densities across different initial densities in the sparse graph (“G2 circuit” test case).
Density (D) $\kappa(L_{G},L_{H})$ GRASS-D inGRASS-D
12.7% $\rightarrow$ 32% 56 $\rightarrow$ 190 13.1% 14.8%
11.8% $\rightarrow$ 32% 66 $\rightarrow$ 196 12.0% 13.8%
9.0% $\rightarrow$ 32% 72 $\rightarrow$ 283 11.0% 11.4%
7.6% $\rightarrow$ 32% 87 $\rightarrow$ 373 8.6% 8.8%
6.6% $\rightarrow$ 32% 103 $\rightarrow$ 432 7.7% 7.9%
IV-C Runtime Scalability of inGRASS
Figure 4: Runtime scalability comparison between GRASS and inGRASS.
Fig. 4 illustrates the runtime scalability of inGRASS compared with GRASS on a logarithmic scale for various test cases. It is demonstrated that inGRASS is faster than GRASS by over 200×200$\times$200 × during $10$-iterative updates. The runtime of inGRASS is also shown by adding the one-time setup time to provide a more comprehensive insight into the overall runtime of inGRASS compared with GRASS.
V Conclusion
This work presents inGRASS, a highly scalable and parallel-friendly algorithm for incremental spectral sparsification of large undirected graphs. The algorithm leverages a low-resistance-diameter decomposition (LRD) scheme to decompose the sparsifier into small clusters with bounded effective-resistance diameters. A multilevel resistance embedding framework is introduced for efficiently identifying spectrally-critical edges as well as detecting redundant ones. The proposed inGRASS achieves state-of-the-art results for incremental spectral sparsification of various networks derived from circuit simulations, finite element analysis, and social networks.